use substitution to find all points of intersection of the graphs of x-y=2 and 6x-y=16
Answer by sofiyac(983)

x-y=2
6x-y=16
we will solve the first equations for x by adding y to each side
x=2+y next plug that into second equation ***
6(2+y)-y=16
12+6y-y=16
12+5y=16
5y=4
y=4/5 next plug that into equation marked with ***
x=2+4/5=10/5+4/5=14/5
(14/5,4/5)